### v1

# The New LLM Bottleneck: A Systems Perspective on Latent Attention and Mixture-of-Experts

Abstract
Computational workloads composing traditional Transformer models are starkly bifurcated. Multi-Head Attention (MHA) is memory-bound, with low arithmetic intensity, while feedforward layers are compute-bound. This dichotomy has long motivated research into specialized hardware to mitigate the MHA bottleneck.
This paper argues that recent architectural shifts, namely Multi-head Latent Attention (MLA) and Mixture-of-Experts (MoE), challenge the premise of specialized attention hardware. We make two key observations. First, the arithmetic intensity of MLA is over two orders of magnitude greater than that of MHA, shifting it close to a compute-bound regime well-suited for modern accelerators like GPUs. Second, by distributing MoE experts across a pool of accelerators, their arithmetic intensity can be tuned through batching to match that of the dense layers, creating a more balanced computational profile.
These findings reveal a diminishing need for specialized attention hardware. The central challenge for next-generation Transformers is no longer accelerating a single memory-bound layer. Instead, the focus must shift to designing balanced systems with sufficient compute, memory capacity, memory bandwidth, and high-bandwidth interconnects to manage the diverse demands of large-scale models.
I Introduction
Transformer-based large language models (LLMs) [52] have achieved remarkable accuracy across a variety of natural language processing tasks [11]. An LLM summarizes/prefills an input (i.e., a sequence of tokens); then, it generates an output token once per decode step. A conventional LLM consists of a sequence of decoder blocks, each comprising a Multi-Head Attention (MHA) block and a Feedforward Network block. To improve scalability and efficiency, recent architectures have incorporated optimizations such as Multi-head Latent Attention (MLA) [11], which first appears at DeepSeek-V3 [27] and reduces the memory footprint of attention, and Mixture-of-Experts (MoE) [9, 15, 29, 55, 22], which enhances model capacity without a proportionally rise in compute cost by activating only a subset of experts per token [47].
Maximizing accelerator utilization when serving these models is critical for improving throughput and end-to-end latency [1]. Serving LLMs in production requires deploying systems that integrate hundreds or even thousands of accelerators (e.g., GPUs [33] and TPUs [50]) to handle high query volumes and large models [4, 41]. Inefficient utilization results in compute resources remaining idle, higher infrastructure costs, and failure to meet service-level objectives (SLOs) [59].
A key factor in serving LLMs efficiently is arithmetic intensity (a.i.), the ratio of arithmetic operations to memory access, measured in operations per byte (Op/B). The ridge point of an accelerator, a concept from the roofline model [54], defines the a.i. at which performance transitions from being memory-bound to compute-bound. To fully exploit an accelerator’s capabilities, the a.i. of each model layer should be configured to approach this ridge point.
In this paper, we argue that MLA and MoE fundamentally reshape the computational landscape of LLM inference. Crucially, both techniques significantly reduce per-request memory demands, especially in the decode stage, thereby enabling much larger batch sizes during inference. By introducing a latent space to attention, MLA significantly reduces the KV\$ size, which is a major bottleneck in conventional LLMs using MHA. Small KV\$ size enables much larger bath sizes in the decode stage without exceeding the memory capacity of the accelerators. Furthermore, by applying layer reordering, the a.i. of core attention in MLA is shifted toward the ridge point of modern accelerators while reducing memory access, enabling significantly higher throughput with large batch sizes while still satisfying service-level objectives (SLOs).
Mixture-of-Experts (MoE) enhances LLM scalability by sparsely activating a subset of experts for each token, enabling larger model capacity without proportional increases in computation. When combined with the increased batch sizes enabled by MLA, MoE can achieves a.i. near to the ridge point of the acceleratos in its expert FC layers with huge batch sizes. We analyze the required batch size to match the Op/B of FC layers in MoE blocks, also attention blocks, while considering the two primary factors, the memory capacity and SLO, which limiting the batch size.
Finally, we present an end-to-end analysis of LLM serving that synthesizes the insights from MLA and MoE. We demonstrate how these architectural innovations synergistically increase the feasible batch size, allowing the system to operate near the accelerator’s ridge point. Furthermore, we highlight the critical role of high-bandwidth interconnects in mitigating communication overheads across accelerators. Our evaluation shows that high bandwidth interconnects, such as NVLink, are essential for minimizing latency in expert dispatch and aggregation, enabling systems to meet service-level objectives (SLOs) with fewer devices. Together, these insights offer a guideline for designing high-throughput, low-latency LLM serving systems.
The key contributions of this paper are as follows:
  • We show that layer reordering in MLA increases the arithmetic intensity of core attention, enabling large batch sizes and higher throughput near the accelerator’s ridge point.
  • We analyze MoE execution and identify the batch size required to match the Op/B of expert FC layers, considering memory and latency constraints.
  • We present an end-to-end analysis of LLM serving systems, highlighting how MLA, MoE, and interconnect bandwidth together enable efficient, scalable inference.
II Background
II-A Standard LLM architecture and its layers
Despite the rapid advancements in LLM algorithms [19, 30], the transformer-based architectures [52], especially the ones using decoders only [7], remain the standard backbone across most modern LLMs (Figure 1). A transformer consists of a series of $n_{\mathrm{dec}}$ decoder blocks. Given an input sequence of length $\ell$ constituting an inference (serving) request, each token (e.g., word) is embedded into a $d_{\mathrm{emb}}$-dimensional hidden state, forming a hidden state matrix $\mathbf{H}_{\ell}\in\mathbb{R}^{\ell\times d_{\mathrm{emb}}}$, then provided as input to the decoder. The matrix passes through a series of decoder blocks, each equipped with its own trained weight parameters, and is transformed into a single output vector.
LLM inference consists of prefill (summarization) stage and decode (generation) stage. The prefill stage processes the entire input hidden stage matrix ($\ell=L_{\mathrm{in}}$) to generate the first output token. Subsequently, the decode stage generates the remaining tokens auto-regressively, where each step takes the single, previously generated token ($\ell=1$) as input to produce the next one, continuing until an end-of-sequence token is generated or the output sequence reach the maximum output token.
Figure 1: Transformer-decoder-based LLM architecture.
A decoder block in a standard LLM consists of two blocks: Multi-Head Attention (MHA) block and Feed-Forward Network (FFN) block. In MHA, a hidden state matrix $\mathbf{H}$ is first linearly transformed (projected) into Query (Q), Key (K), and Value (V) matrices by passing through fully connected (FC) layers with pre-trained weights, which are then split into $n_{\mathrm{hd}}$ heads. Eq. 1 shows how $\mathbf{Q},\mathbf{K}$, and $\mathbf{V}$ for each head is computed where $L$ denotes the current sequence length, defined as the sum of the input sequence length $L_{in}$ and the number of tokens generated so far. During auto-regressive decoding, a KV cache (KV\$) stores past K and V values to maintain contextual information without costly recomputation. Consequently, only the new K and V vectors for the current input token are computed and appended to this cache.
$\begin{split}\underbrace{\mathbf{Q}_{i}}_{\mathbb{R}^{\ell\times\frac{d_{\mathrm{dec}}}{n_{\mathrm{hd}}}}}=\underbrace{\mathbf{H}_{\ell}}_{\mathbb{R}^{\ell\times d_{\text{emb}}}}\cdot\underbrace{\mathbf{W}_{\mathbf{Q}_{i}}}_{\mathbb{R}^{d_{\text{emb}}\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}\\ \underbrace{\mathbf{X}_{i}}_{\mathbb{R}^{L\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}=\underbrace{\mathbf{H}_{L}}_{\mathbb{R}^{L\times d_{\text{emb}}}}\cdot\underbrace{\mathbf{W}_{\mathbf{X}_{i}}}_{\mathbb{R}^{d_{\text{emb}}\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}\\ \text{for }\mathbf{X}\in\{\mathbf{K},\mathbf{V}\}\text{ and }i\in[1,n_{\text{hd}}]\end{split}$
Each head performs a sequence of operations, referred to as core-attention, independently. Core-attention computes score (Eq. 2), softmax, and context (Eq. 3) operations.
$\begin{split}\underbrace{\mathbf{S}_{i}}_{\mathbb{R}^{\ell\times(L)}}=\underbrace{\mathbf{Q}_{i}}_{\mathbb{R}^{\ell\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}\cdot\underbrace{[\mathbf{K}_{i}]^{\text{T}}}_{\mathbb{R}^{\frac{d_{\text{dec}}}{n_{\text{hd}}}\times L}}\end{split}$
$\underbrace{\mathbf{O}_{i}}_{\mathbb{R}^{\ell\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}=\underbrace{\text{Softmax}({\frac{\mathbf{S}_{i}}{\sqrt{d_{\text{dec}}/{n_{\text{hd}}}}}})}_{\mathbb{R}^{\ell\times L}}\cdot\underbrace{[\mathbf{V}_{i}]}_{\mathbf{R}^{L\times\frac{d_{\text{dec}}}{n_{\text{hd}}}}}$
Table I: Arithmetic performance, main-memory bandwidth and capacity of deep-learning accelerators
V100 SXM2 [31] A100 SXM4 [32] H200 SXM5 [34] B200 SXM6 [36] TPU V5P [51] TPU V7 [50] MI325X [3]
BF16 performance (TFLOPS) 125 312 989.5 2250 459 2307 1307.4
Memory bandwidth (GB/s) 900 2039 4800 8000 2765 7400 6000
HBM capacity per GPU (GB) 32 80 141 192 95 192 256
Ridge point (BF16) 138.89 153.02 206.15 281.25 166 320.42 217.9
Lastly, MHA performs an FC layer, attention output projection, generating MHA’s output U.
An FFN block in recent LLMs consists of three FC layers and one nonlinear activation. Until GPT-3 [21], most models used an FFN block with only two FC layers and one activation. However, modern LLMs typically employ three FC layers, which improves response quality at the cost of additional computation by introducing gated activation functions [46].
Rotary Positional Embedding (RoPE) [49] is common in modern LLMs to inject positional information into token generations. RoPE encodes each token’s position through a unique rotational transformation based on its position index applied to $\mathbf{Q}$ and $\mathbf{K}$ before the core attention layer while leaving $\mathbf{V}$ unchanged.
For the FC layers, all requests share the same weights in LLM inference. Batching multiple requests enables the FC weight reuse, reducing memory access overhead. As memory bandwidth scales more slowly than compute capability in modern accelerators (Table I), it becomes a major bottleneck, especially in the decode stage, where GEMV operations offer limited data reuse. While the batching improves throughput, it does not reduce per-request latency, and the batch size $B$ is bounded by SLOs [59]. Moreover, as each request maintains its own KV\$, a larger $B$ increases memory capacity requirements, making memory capacity another key limiter for the maximum feasible value.
II-B Representative optimizations for LLM
To efficiently serve LLMs, a myriad of optimizations—spanning both algorithmic strategies and hardware-aware techniques—have been proposed in response to ever-growing demand for LLMs.
Grouped-Query Attention (GQA): During the decode stages ($\ell=1$), the core attention of MHA performs multiple GEMVs (see Eq. 2 and Eq. 3) between the query vectors and request-specific KV\$ matrices. These layers are highly memory-bound with a fixed arithmetic intensity (a.i.) of 1; thus, they make poor use of the accelerator’s compute resources [8, 20].
GQA [2] mitigates this issue. In GQA, a head of KV\$ is shared by multiple Q heads, reducing the total KV\$ size by the number of queries in a group. Reusing a KV\$ across multiple heads, a.i. improves while deteriorating model accuracy. An extreme variant, Multi-Query Attention (MQA) [45], enforces all Q heads to share a single KV head. However, due to the significant degradation in service quality observed with MQA compared to MHA and GQA, we will not discuss it further.
[…]
$\underbrace{\mathbf{C}_{\mathrm{Q}}}_{\mathbb{R}^{\ell\times d_{\mathrm{Qco}}}}=\mathbf{H}_{\ell}\cdot\underbrace{\mathbf{W}_{\mathrm{CQ}}}_{\mathbb{R}^{d_{\mathrm{emb}}\times d_{\mathrm{Qco}}}},\underbrace{\mathbf{C}_{\mathrm{KV}}}_{\mathbb{R}^{L\times d_{\mathrm{KVco}}}}=\mathbf{H}_{L}\cdot\underbrace{\mathbf{W}_{\mathrm{CKV}}}_{\mathbb{R}^{d_{\mathrm{emb}}\times d_{\mathrm{KVco}}}}$
$\mathbf{S}_{i}=\mathbf{Q}_{i}\cdot(\mathbf{K}_{i})^{\text{T}}\!=\!\underbrace{(\mathbf{C}_{\mathrm{Q}}\cdot\mathbf{W}_{\mathrm{DQ}_{i}})}_{\mathclap{\mathbb{R}^{\ell\times\frac{d_{\mathrm{dec}}}{n_{\mathrm{hd}}}}}}\cdot(\underbrace{\mathbf{C}_{\mathrm{KV}}}_{\mathbb{R}^{L\times d_{\mathrm{KVco}}}}\cdot\underbrace{\mathbf{W}_{\mathrm{DK}_{i}}}_{\mathclap{\mathbb{R}^{d_{\mathrm{KVco}}\times\frac{d_{\mathrm{dec}}}{n_{\mathrm{hd}}}}}})^{\text{T}}&\end{aligned}$
$\mathbf{O}_{i}=\mathrm{Softmax}({\frac{\mathbf{S}_{i}}{\sqrt{d_{\mathrm{dec}}/{n_{\mathrm{hd}}}}}})\cdot(\mathbf{C}_{\mathrm{KV}}\cdot\underbrace{\mathbf{W}_{\mathrm{DV}_{i}}}_{\mathclap{\mathbb{R}^{d_{\mathrm{KVco}}\times\frac{d_{\mathrm{dec}}}{n_{\mathrm{hd}}}}}})$
Notably, instead of retaining the separate $n_{\mathrm{hd}}$ heads of key and value tensors, $\mathbf{K}_{i},\mathbf{V}_{i}\in\mathbb{R}^{L\times\frac{d_{\mathrm{dec}}}{n_{\mathrm{hd}}}}$, MLA stores a single latent KV\$, $\mathbf{C}_{\mathrm{KV}}\in\mathbb{R}^{L\times d_{\mathrm{KVco}}}$. It is then used to reconstruct both the key and value heads. As the latent KV\$ dimension $d_{\mathrm{KVco}}$ is much (usually more than an order of magnitude) smaller than both $d_{emb}$ and $d_{dec}$, it significantly reduces memory usage (by a factor of $\frac{2d_{dec}}{d_{\mathrm{KVco}}}=64$ for DeepSeek-R1) and decreases data movement during the core-attention. However, these gains come at the cost of additional overhead for on-demand KV decompression per decode stage.
[…]
$\begin{split}\mathbf{S}_{i}&=\mathbf{Q}_{i}\cdot(\mathbf{C}_{\mathrm{KV}}\cdot\mathbf{W}_{\mathrm{DK}_{i}})^{\text{T}}\\ &=\mathbf{Q}_{i}\cdot({\mathbf{W}_{\mathrm{DK}_{i}}}^{\mathrm{T}}\cdot{\mathbf{C}_{\mathrm{KV}}}^{\text{T}})\\ &=(\mathbf{Q}_{i}\cdot{\mathbf{W}_{\mathrm{DK}_{i}}}^{\text{T}})\cdot{\mathbf{C}_{\mathrm{KV}}}^{\text{T}}\\ \mathbf{S}&=\begin{bmatrix}\mathbf{Q}_{1}\cdot{\mathbf{W}_{\mathrm{DK}_{1}}}^{\text{T}}\\ \mathbf{Q}_{2}\cdot{\mathbf{W}_{\mathrm{DK}_{2}}}^{\text{T}}\\ \vdots\\ \mathbf{Q}_{n_{\mathrm{hd}}}\cdot{\mathbf{W}_{\mathrm{DK}_{n_{\mathrm{hd}}}}}^{\mathbf{T}}\end{bmatrix}\cdot{\mathbf{C}_{\mathrm{KV}}}^{\text{T}}\end{split}$
A similar reordering can also be applied to the context layer:
$\begin{split}\mathbf{O}_{i}&=\text{Softmax}({\frac{\mathbf{S}_{i}}{\sqrt{d_{\mathrm{dec}}/{n_{\mathrm{hd}}}}}})\cdot(\mathbf{C}_{\mathrm{KV}}\cdot\mathbf{W}_{\mathrm{DV}_{i}})\\ &=\!(\text{Softmax}({\frac{\mathbf{S}_{i}}{\sqrt{d_{\mathrm{dec}}/{n_{\mathrm{hd}}}}}})\cdot\mathbf{C}_{\mathrm{KV}})\cdot\mathbf{W}_{\mathrm{DV}_{i}}\end{split}$
The reordering transforms the score and context layers from GEMV to GEMM operations without affecting their results. Combined with the use of latent space, this reordering synergistically improves the a.i. and reduces the memory accesses of core-attention (further elaborated in §V).
Mixture of Experts (MoE): Although a common belief is that larger models with more parameters produce higher-quality responses [24], the substantial computational overhead associated with scaling LLMs hinders their adoption. MoE [15] addresses this issue by employing sparse activation in FFN blocks called an expert. MoE introduces a pool of experts and activates only a small subset of these experts for each input. In particular, recent LLMs [12, 11, 29] adopt a hybrid architecture consisting of two types of experts: a shared expert and routed experts. The former is activated for every token during inference, whereas the latter are selectively activated based on a routing mechanism that dynamically assigns $n_{\mathrm{k}}$ experts out of $n_{\mathrm{e}}$ experts to each token. The computational procedure of an MoE block can be described as
$\begin{split}\text{Router}(\mathbf{u})\!=\!\mathbf{E}\in\{1,2,\cdots,n_{e}\},\>|\mathbf{E}|=n_{k}<n_{e}\\ \text{MoE}(\mathbf{u})\!=\!(\sum_{e\in\mathbf{E}}\text{Expert}_{e}(\mathbf{u}))\!+\!\text{Expert}_{\mathrm{shared}}(\mathbf{u})\end{split}$
By utilizing only $n_{k}$ experts ($n_{k}<n_{e}$), along with one shared expert per token at runtime, MoE scales the model with significantly lower computational overhead. The values of $n_{k}$ and $n_{e}$ vary across different models: DeepSeek-R1 [11] employs eight routed experts selected from a pool of 256, while Llama 4 Maverick [29] uses a single routed expert selected from a pool of 128.
II-C Hardware efficiency & arithmetic intensity (a.i.)
When serving LLMs, three key factors determine service latency and throughput: arithmetic performance, memory bandwidth, and memory capacity of an accelerator. High arithmetic performance enables fast execution of compute-intensive layers, such as GEMMs on large tensors during the prefill stage (e.g., Eq. 1). To fully leverage this arithmetic capability, sufficient memory bandwidth is essential to quickly supply the data required for these computations. Such high bandwidth can only be fully utilized when the memory capacity is large enough to accommodate the entire working set; otherwise, data must be fetched from lower-bandwidth sources (e.g., via a PCIe link), severely limiting performance.
[…]
II-D LLM serving systems
LLMs have progressively scaled up in size to improve quality, their computation and memory demands now far exceed the capacity of a single accelerator. For example, DeepSeek-R1 requires over 1250 GB of memory with BF16 precision, far exceeding the main memory capacity of the latest accelerators (e.g., 192 GB per NVIDIA B200 GPU, see Table I). Even with sufficient memory capacity, limited compute resources may still lead to unacceptable latency or low throughput. Thus, LLM serving systems have evolved to utilize multiple accelerator instances.
In LLM serving systems, various parallelisms are employed to efficiently distribute the model weights and computations across accelerators [14, 6]. The most prominent approaches include data parallelism (DP), tensor parallelism (TP[48], and expert parallelism (EP[15, 43]. DP replicates the model weights across multiple accelerators and splits the input batch into sub-batches, processing each sub-batch independently on a separate accelerator without inter-accelerator communication. However, it introduces memory inefficiency as identical model weights are stored repeatedly on each accelerator.
TP partitions activation, weights, or both between accelerators, allowing each to compute partial results of a single operation in parallel. It improves the utilization of memory and compute resources, but introduces inter-accelerator communication as partial results must be exchanged and aggregated during execution.
EP distributes experts in MoE blocks evenly across accelerators. The output vector of an attention block is dispatched to the appropriate accelerators that store the routed experts. After the corresponding MoE blocks have been computed, the results in different accelerators are combined. This dispatching and combining incur additional communication overhead. $deg_{\mathrm{TP}}$, $deg_{\mathrm{DP}}$, and $deg_{\mathrm{EP}}$ denote the degree of TP, DP, and EP, respectively. Moreover, different types of parallelism can be applied selectively at the block level, enabling optimization based on the computational characteristics of each block.
Pipeline parallelism (PP), an inter-layer model parallelization technique, partitions sequential layers into multiple accelerators to address memory capacity limitations with lower communication cost over TP. However, TP is often preferred over PP in LLM inference as it directly reduces the latency of computations, whereas PP can not. Another core weakness of PP is the creation of inevitable bubbles, where accelerators sit idle as they have to wait for the input data from the previous layer computed by other accelerators in the pipeline. Although it can be mitigated by processing several micro-batches to keep the pipeline full, this method’s effectiveness diminishes with smaller batch sizes, leading to hardware under-utilization. Thus, in this paper, we focus on TP rather than PP.
As the LLM serving system gets bigger, interconnection becomes a critical factor in determining system performance and scalability. Efficient interconnect bandwidth topology directly impacts the cost and latency of communication between accelerators [10, 35, 12].
Table II: Symbols used throughout this paper, their descriptions, and the exemplar parameters used in DeepSeek-R1 [11]
Term Description DeepSeek-R1
$d_{\mathrm{emb}}$ Embedding dimension (dim) 7168
$n_{\mathrm{hd}}$ Number of heads 128
$d_{\mathrm{hd}}$ Head dim 128
$d_{\mathrm{dec}}$ Decompressed Q, K, V dim ($=\!n_{\mathrm{hd}}\!\times\!d_{\mathrm{hd}}$) 16384
$d_{\mathrm{Qco}},d_{\mathrm{KVco}}$ Compressed Q, KV dim 1536, 512
$d_{\mathrm{RoPE}}$ Rotary PE dim 64
$d_{\mathrm{MoE}}$ MoE intermediate dim 2048
$n_{\mathrm{e}}$ Number of routed experts 256
$n_{\mathrm{k}}$ Number of routed experts per token 8
III Experimental Setup
To evaluate LLM serving performance in various configurations, we built an in-house simulator based on the LLMSimulator [44], which was used in [56]. Our in-house simulator is configured to model the latest NVIDIA B200 GPU, whose key parameters are specified in Table I. By default, we assume all GPUs in a group are fully connected via NVLink fifth generation, providing 900GB/s of unidirectional bandwidth following the NVL72 system topology [33]. For each experiment, we specify the number of GPUs per group and note when InfiniBand XDR (100 GB/s) is used for inter-group communication. All experiments used DeepSeek-R1 (key parameters summarized in Table II and specified in Table IV) with BF16 data size for all parameters, KV\$, and activations. Following [40, 59, 35, 16], the system is disaggregated, with the prefill and decode stages executed on separate machines.
We also varied the batch size by stage. In the decode stage, where a single vector is processed for each request, batching is crucial for achieving high hardware utilization, while the prefill stage derives less benefit from batching as it already processes a large sequence of tokens, which can effectively saturate the compute resources [59, 1, 13]. Since the decode stage dominates overall execution time [39], our analysis primarily focuses on the decode stage.
IV Demystifying the Compound Interplay of Algorithm–Hardware–Parallelism
Figure 3: Time per output token and per-device throughput of DeepSeek-R1 and GPT-3 across varying sequence length $L$ and batch size $B$. The blurred area indicates configurations where out-of-memory errors occurred and the red dashed line represents the time required to read all data from HBM, which determines the maximum feasible combinations of $(B,L)$. The experiment assumes a 32 NVIDIA B200 GPU system.
IV-A The latest LLM wears MLA and MoE
Contemporary LLMs, leveraging MLA and MoE, demonstrate significant enhancements in both inference latency and throughput. Figure 3 illustrates this by comparing the time per output token (TPOT) and token generation throughput between DeepSeek-R1 and GPT-3 during inference on a 32 B200 GPU system. Hereafter, the blurred area indicates configurations where out-of-memory errors occurred. This comparison reveals an impressive throughput improvement of up to 53.67×53.67$\times$53.67 ×, alongside a notable increase in the maximum feasible batch size ($B$). We delve into how MLA and MoE impact the computational characteristics of primary computing blocks within LLM execution on typical multi-accelerator serving systems. Given the alterations in model structure and weight distribution across blocks, we reassess optimal serving strategies based on our observations. We especially focus on long sequences (over 1024) in this paper, following recent trends, such as chain of thought [58, 53], that keep increasing the average sequence lengths [38, 5, 25]. Our observations are highlighted in separate boxes and categorized into two types: primary (Observation-P) and supportive (Observation-S).
IV-B limitations of conventional LLMs
Core attention’s memory and computational demands remain consistent per request, even with batching. This is because, unlike FC layers that benefit from batching for weight memory access amortization, core attention processes activations unique to each request. Thus, its a.i. is not altered by batching, as mandated by the model architecture.
Observation-S1: The a.i. of core attention is decided by the model architecture, regardless of the batch size.
Figure 4 shows total model weight size breakdowns for a conventional LLM (GPT-3): Until FC layers batch enough requests to reach the ridge point $RP_{\mathrm{acc}}$, the runtime of each layer is proportional to the memory usage, as they are all bounded by reading data from memory.
As the sequence length and batch size ($B$) increase, core attention becomes the main bottleneck in the decode stage, limiting throughput and leaving hardware underutilized [39, 56, 20]. The maximum feasible $B$ in an LLM system is determined by the memory capacity that remains after storing the model parameters because the residual capacity accommodates the per-request KV\$. FC layers that use weights are in a memory-bound region when $B$ is smaller than $RP_{\mathrm{acc}}$. With an insufficient maximum $B$, time per output token (TPOT) is bounded by the time required to read the entire memory, leaving computing resources in an accelerator idle.
Although larger models generally translate to higher quality outputs, a dense model must touch every parameter at inference time, making the computational cost grow linearly with model size. Figure 4 compares the number of activated parameters per token between DeepSeek-R1, which employs a MoE network, and GPT-3 that uses a dense FFN. While the model size of GPT-3 is much smaller than that of DeepSeek-R1, the amount of computation per token didn’t scale accordingly.
Figure 4: Memory usage comparison of DeepSeek-R1 and GPT-3, including attention weight, FFN/MoE weight, and the KV\$ for 256K tokens with BF16 precision. While GPT-3 needs to use full model parameters during inference, DeepSeek-R1, which employs MoE architecture, activates only 37B parameters per token, requiring approximately 70GB of memory usage.
IV-C Insights into contemporary accelerators for LLMs
An algorithm’s performance must be evaluated relative to the hardware’s ridge point. For example, in an FC layer, increasing batch size raises a.i. and throughput prior to reaching this point. Beyond the ridge point, the system becomes compute-bound, and batching further requests only increases latency without yielding additional throughput gains. Thus, the implications of changing an algorithm’s a.i. (Op/B) are meaningful when interpreted in the context of this ridge point.
The B200 GPU provides approximately 18$\times$ higher arithmetic throughput compared to the V100 GPU, representing the most significant improvement among the accelerators listed in Table I. However, because both arithmetic performance and memory bandwidth scale with technology, the B200’s ridge point increases by only a modest factor of 2.02 compared to the V100. Thus, contemporary accelerators exhibit ridge points within a narrow range of 200-400 Op/B.
Observation-S2: Modern accelerators typically have the ridge points ($\mathrm{RP_{acc}}$) in the range of 200 to 400 Op/B.
V Insights on Attention
Table III: Comparison of MLA with and without the reordering in terms of FLOPS, memory access, and a.i. for prefill and decode stages.
Layer Phase Reordering FLOPS Asymptotic Memory Access a.i. a.i. in DeepSeek-R1
Prefill K decompress without $B2Ld_{\mathrm{KVco}}n_{\mathrm{hd}}d_{\mathrm{hd}}$ $2BLn_{\mathrm{hd}}d_{\mathrm{hd}}$ $\approx d_{\mathrm{KVco}}$ $\approx 512$
with $B2Ld_{\mathrm{KVco}}n_{\mathrm{hd}}d_{\mathrm{hd}}$ $2B(Ln_{\mathrm{hd}}d_{\mathrm{hd}}+Ln_{\mathrm{hd}}d_{\mathrm{KVco}})$ $\approx(d_{\mathrm{hd}}^{-1}+d_{\mathrm{KVco}}^{-1})^{-1}$ $\approx 100$
Score without $B2n_{\mathrm{hd}}L^{2}d_{\mathrm{hd}}$ $2Bn_{\mathrm{hd}}L^{2}$ $\approx d_{\mathrm{hd}}$ $\approx 128$
with $B2n_{\mathrm{hd}}L^{2}d_{\mathrm{KVco}}$ $2Bn_{\mathrm{hd}}L^{2}$ $\approx d_{\mathrm{KVco}}$ $\approx 512$
Decode K decompress without $B2d_{\mathrm{KVco}}Ln_{\mathrm{hd}}d_{\mathrm{hd}}$ $2BLn_{\mathrm{hd}}d_{\mathrm{hd}}$ $\approx d_{\mathrm{KVco}}$ $\approx 512$
with $B2d_{\mathrm{KVco}}n_{\mathrm{hd}}d_{\mathrm{hd}}$ $2Bd_{\mathrm{KVco}}n_{\mathrm{hd}}$ $\approx d_{\mathrm{hd}}$ $\approx 128$
Score without $B2n_{\mathrm{hd}}Ld_{\mathrm{hd}}$ $2Bn_{\mathrm{hd}}d_{\mathrm{hd}}L$ $\approx 1$ $\approx 1$
with $B2n_{\mathrm{hd}}Ld_{\mathrm{KVco}}$ $2B(d_{\mathrm{KVco}}L+n_{\mathrm{hd}}L)$ $\approx(d_{\mathrm{hd}}^{-1}+d_{\mathrm{KVco}}^{-1})^{-1}$ $\approx 100$
We analyze the computational characteristics of a key computation block, MLA, which is composed of multiple FC layers and a core attention.
Figure 5: (a) Normalized latency of the attention block in the decode stage without reordering compared to a reordered Attention block. (b) and (c) shows execution time ratio of each layer in the attention block in the decode stage with and without reordering, across varying sequence length $L$ and batch size $B$. The experiments assume a 32 NVIDIA B200 GPU system.
V-A Introducing latent space to attention
MLA employs low-rank joint compression to the attention block, reducing its projection weight size and the associated FC layer compute. Thus, attention block’s parameter footprint shrinks substantially both in absolute size and as a fraction of total model parameters. With this reduced weight burden, replicating its FC layer parameters across devices is far more tractable, making DP for the attention block more feasible.
[…]
V-B Impact of Reordering MLA
The reordering explained in §II-B drastically reduces the latency of attention blocks during the decode stage, whereas it rather increases the latency during the prefill stage. It is because the reordering in MLA significantly changes computation characteristics, such as FLOPs, memory access, and a.i..
[…]
Without reordering, the core attention in MLA preserves the same computational flow as MHA except for the runtime compression and decompression steps immediately preceding it. The required computation (FLOPs) and memory access for the K decompression and score layers are proportional to $B$ and $L$ during the decode stage (see Table III), whereas those of the FC layers do not scale with $L$. Thus, KV decompression and core attention dominate the execution time of an attention block as $L$ increases (see Figure 5(b)). At $B\!=\!256$ and $L\!=\!4096$, KV decompression and core attention account for 59% and 40% of an attention block latency, respectively.
Observation-P1: In the decode stages of MLA without reordering, KV decompression and core attention layers dominate the runtime of MLA’s attention blocks.
After applying the layer reordering, MLA multiplies $\mathbf{W}_{\mathrm{DK}}$ with $\mathbf{Q}$. As the size of $\mathbf{Q}$ is independent of $L$ in the decode stage, reordering substantially reduces computation in a K decompression step by a factor of $L$, as this layer no longer decompresses an entire $\mathbf{C}_{\mathrm{KV}}$ (Table III). The portion of K decompression—previously the dominant component—has been significantly reduced due to reordering (see Figure 5(c)).
By contrast, this benefit disappears in the prefill stage as the size of $\mathbf{Q}$ is proportional to $L\!=\!L_{\mathrm{in}}$, making the computational cost of K decompression unchanged. The reordering rather increases the amount of computation required for the score layer by $\nicefrac{{d_{\mathrm{KVco}}}}{{d_{\mathrm{hd}}}}$ (4 for DeepSeek-R1) times in both stages, as it replaces the original score layer with a multiplication between $\mathbf{Q}_{i}\cdot{\mathbf{W}_{\mathrm{DK_{i}}}}^{\mathrm{T}}$ and ${\mathbf{C}_{\mathrm{KV}}}^{\mathrm{T}}$ (see Eq. 7), where one of the matrix dimensions becomes $d_{\mathrm{KVco}}$ instead of $d_{\mathrm{hd}}$.
Observation-P2: While the layer reordering maintains or reduces the FLOPs of KV decompression, it instead increases the FLOPs of core attention in both stages.
[…]
FlashMLA [23], a GPU-optimized implementation, further doubles this Op/B by reusing ${\mathbf{C}_{\mathrm{KV}}}$ loaded during the score layer in the subsequent context layer. A $2n_{\mathrm{hd}}$ Op/B (i.e., 256 in DeepSeek-R1) closely approaches the ridge point of modern accelerators (Obs.S2). Since the a.i. of core attention is near $RP_{\mathrm{acc}}$ of modern accelerators, its performance is balanced between compute-bound and memory-bound. This means the time spent on computation is still approximately equivalent to the time spent on $\mathbf{C}_{\mathrm{KV}}$ access, thereby the throughput of core attention improved approximately by $2d_{\mathrm{dec}}/d_{\mathrm{KVco}}$ ($=\!64$ for DeepSeek-R1).
In contrast, reordering increases the latency of attention blocks during the prefill stage. In the KV decompression layer, the FLOPs remain unchanged, while memory access increases, leading to longer execution time. For the core attention, memory access and FLOPs grow when the layer reordering is applied because its memory footprint is dominated by input/output activations of core attention, unlike the decode stages. Specifically, layer reordering expands the activation’s dimension per head from $d_{\mathrm{hd}}$ to $d_{\mathrm{KVco}}$. Thus, the additional computation outweighs any potential gains. Applying the layer reordering deteriorates the performance of attention blocks during the prefill stage (Obs.P2).
Putting it all together, layer reordering significantly reduces the latency of KV decompression and core attention within the overall attention block latency in the decode stage (see Figure 5(a)). Thus, layer reordering improves the execution time of the attention block in the decode stage by 119.07$\times$, when $B\!=\!256$ and $L\!=\!4096$. When $B\!=\!1$ and $L\!=\!4096$, the attention block’s latency in the prefill stage with reordering is 2.02$\times$ higher than that without reordering. Hereafter, the prefill stage uses MLA without reordering and the decode stage uses MLA with reordering.
Observation-P3: Layer reordering enhances the attention block’s throughput by substantially improving its a.i., primarily by reducing memory access despite the additional computational overhead in the decode stage; however, it degrades performance in the prefill stage.
MLA diminishes the need for specialized attention hardware. In contrast to MHA or GQA, where core attention in decode stages typically shows low a.i. (Op/B in the 1–8 range) [56], prompting the development of processing-in-memory or near-data-processing architectures [39, 56, 20, 28] for acceleration, MLA and reordering significantly increase the Op/B of core attention, approaching $\mathrm{RP_{acc}}$. Consequently, previous attention-specific hardware is less effective for accelerating MLA attention because performance is no longer constrained by memory bandwidth.
V-C Parallelism on MLA
TP offers little latency benefit once reordering is applied in the attention block. Although reordering massively lowers the core attention latency, the latency still scales with $B$ and $L$ in the decode stage and dominates at large $B$ and $L$ (Figure 5(c)). When heads are independent, as in MHA, head-wise TP can distribute the KV\$ across multiple accelerators, reducing latency. Also, a.i. is preserved within a head as it is not affected by TP.
In the reordered MLA, all heads share the same $\mathbf{C}_{\mathrm{KV}}$; thus, the capacity benefit from TP diminishes. As depicted in Figure 6, TP reduces the number of heads batched on each accelerator, thereby reducing the a.i. by ${deg}_{\mathrm{TP}}$.
This Op/B reduction lowers per-accelerator computational throughput by around ${deg}_{\mathrm{TP}}$$\times$, effectively canceling out the performance gain from scaling out to ${deg}_{\mathrm{TP}}$ accelerators. Figure 7 compares the latency impact of TP on a reordered and not-reordered MLA attention block. Although the FC layers in an attention block benefit from TP, the overall latency improvement is limited because the core attention takes the majority of the runtime of the attention block. Moreover, each head is multiplied by full $\mathbf{C}_{\mathrm{KV}}$, requiring all ${deg}_{\mathrm{TP}}$ accelerators to store redundant copies of $\mathbf{C}_{\mathrm{KV}}$, leading to memory inefficiency.
Observation-P4: In the decode stage with reordering, tensor parallelism fails to provide meaningful latency reduction.
Figure 6: Computation flow of an MLA attention block with ${deg}_{\mathrm{TP}}=2$ and $n_{\mathrm{hd}}=4$ on our experimental setup.
Figure 7: Latency of the Attention block with and without reordering as batch sizes and $\mathrm{deg}_{\mathrm{TP}}$ vary when $L\!=\!4096$.
VI Insights on Mixture of Experts
The execution time of MoE blocks is primarily dominated by performing experts and the communication required for dispatching and combining tokens [57]. In a multi-accelerator system, efficient expert computation requires careful design of the deployment strategy, including data distribution and parallelization, to maximize each accelerator’s arithmetic throughput by batching FC layers in each expert, while meeting accelerator memory and SLO constraints. Meanwhile, communication overhead between accelerators in serving system is highly dependent on the system-wide interconnect (e.g., NVLink [17, 36], InfiniBand [18, 37], and optical links [42]) specification. In this section, we analyze the impact of both factors on the performance of MoE blocks.
Figure 8: Per-GPU memory usage for DeepSeek-R1 as more GPUs are used.
VI-A Maximize compute utilization in FC layers
Although both attention and MoE blocks perform FC layers, the effective batch size differs even under a fixed system-wide batch size $B$, due to the distinct parallelism strategies employed in each block. In the attention block, $B$ requests are evenly distributed across $\mathrm{deg_{DP}}$ groups of accelerators, where each group uses TP to process $\frac{B}{\text{deg}_{\text{DP}}}$ requests in parallel. The Op/B of the FC layers in this block is around $\frac{B}{\text{deg}_{\text{DP}}}$. In contrast, the MoE block utilizes EP, which processes each expert on a single accelerator by gathering attention results from other accelerators. Thus, the effective batch size for MoE blocks remains $B$. Considering that each token is routed to $n_{k}$ experts among $n_{e}$ experts, assuming uniform selection [^2] To avoid unbalanced expert load, SwitchTransformer [15] and GShard [26] introduce an auxiliary loss, and DeepSeek-V2 [27] incorporates a bias term to balance the routing among experts. Following these, we assume the expert load is not highly skewed and approximately follows a uniform distribution [56]., each expert processes $B\cdot n_{k}/n_{e}$ tokens. The average a.i. of MoE blocks becomes around $B\cdot n_{k}/n_{e}$. To make the FC layers in both attention and MoE blocks fully utilize target accelerators, the batch size B should satisfy:
$\begin{split}\mathrm{B}\geq\mathrm{B_{attn}}&=\mathrm{RP_{acc}}\cdot\mathrm{deg_{DP}}\\ \mathrm{B}\geq\mathrm{B_{MoE}}&=\mathrm{RP_{acc}}\cdot\frac{n_{e}}{n_{k}}\end{split}$
where $\mathrm{RP_{acc}}$ denotes accelerator’s ridge point and $\mathrm{B_{attn}}$ and $\mathrm{B_{MoE}}$ are the batch sizes that achieve $\mathrm{RP_{acc}}$ in a.i. for the FC layers of an attention and MoE block, respectively. We denote the minimum $B$ that satisfies Eq. 10 as $\mathrm{B_{RP}}=\text{max}(\mathrm{B_{attn}},\mathrm{B_{MoE}})$.
We observe that $\mathrm{B_{attn}}$ is influenced by $\mathrm{deg_{DP}}$ whereas $\mathrm{B_{MoE}}$ is not. Since $n_{e}$ and $n_{k}$ are model parameters, the a.i. of the FC layers in an MoE block is determined once $B$ and the model are fixed. In other words, the batch size that fully utilizes the MoE block, denoted $\mathrm{B_{MoE}}$, is determined solely by the model and the target accelerator’s ridge point, independently of $n_{\mathrm{acc}}$ and the chosen parallelism strategy.
Observation-P5: $B_{\mathrm{MoE}}$ is determined once the model and the target accelerator are fixed.
Figure 9: Throughput-latency graph for the decode stages of GPT-3 and DeepSeek-R1. We assume a 32 B200 GPU system.
VI-B Two primary factors limiting batch size
While batching $\mathrm{B_{RP}}$ requests is desirable, feasible batch size is limited by two factors; memory capacity and SLO.
Memory capacity: To fully utilize the accelerator’s computational resources, data must be served at high bandwidth, which requires the entire working set to reside in the main memory (e.g., HBM) of the accelerators. This working set includes the weights for attention and MoE blocks, as well as KV\$. Since model weights are predetermined, serving systems typically use the remaining memory for activations and KV\$, whose sizes are proportional to $B$. Thus, the memory space requirements for model weights determine the maximum feasible batch size ($\mathrm{B_{cap}}$) as follows:
$\begin{split}\mathrm{B_{cap}}&=\frac{\mathrm{M_{cap}}\cdot n_{\mathrm{acc}}-n_{dec}\cdot(\mathrm{M_{attn}}\cdot{\mathrm{deg_{DP}}}+\mathrm{M_{MoE}})}{{n_{dec}\cdot\mathrm{M_{KVcache}}\cdot L}+\mathrm{M_{act}}(L)}\end{split}$
where $\mathrm{M_{cap}}\cdot n_{\mathrm{acc}}$ denotes the capacity of a system composed with $n_{acc}$ accelerators, each having $\mathrm{M_{cap}}$ capacity. $\mathrm{M_{attn}}$ and $\mathrm{M_{MoE}}$ represent the model weight sizes of a single decoder block’s attention and MoE, respectively. We denote the KV\$ size per token for each decoder block as $\mathrm{M_{KVcache}}$. The activation memory required by a decoder block for a single token on each accelerator, $\mathrm{M_{act}}$, is assumed to depend on the sequence length $L$. As the activation memory is reused across multiple decoder blocks, the term $\mathrm{M_{act}}(L)$ in Eq. 11333The equation assumes that all decoder blocks contain MoE layers, so the total MoE weight stored per accelerator is multiplied by $n_{\mathrm{dec}}$. However, in models like DeepSeek-R1, the first three layers are not MoE layers, and thus the correct factor should be $n_{\mathrm{dec}}-3$. This factor may vary slightly depending on the specific model architecture., does not scale with $n_{\mathrm{dec}}$. To batch $\mathrm{B_{\mathrm{RP}}}$ requests, $\mathrm{B_{cap}}$ should be greater than $B_{\mathrm{RP}}$.
SLO: As excessive batching of requests incurs latency overheads, SLO becomes another limiting factor for feasible batch sizes. In a disaggregated system, the time per output token (TPOT)—a key latency metric in LLM serving—is determined by the latency of each decoding stage and can be expressed as follows:
$\begin{split}\mathrm{TPOT(B,L)}\!&=\!n_{\mathrm{dec}}\cdot\left(\underbrace{\frac{\mathrm{M_{attn}}\cdot{\mathrm{deg_{DP}}}+\mathrm{M_{MoE}}}{n_{\mathrm{acc}}\cdot\mathrm{BW_{Mem}}}}_{\text{model load lat.}}\!+\!\underbrace{\mathrm{T(B,L)}}_{\text{additional lat.}}\right)\end{split}$
where both the first and second terms in the parentheses represent latencies for each decoder block: the first accounts for the latency to read model weights, and the second, $\mathrm{T(B,L)}$, includes additional latency such as memory access time for the KV\$ and activations, communication overhead, and any remaining computation time. The additional latency term is a function of $B$ and $L$.
As memory access time for the KV\$ and activations, along with communication time, is unavoidable while processing each decoder block, the minimum bound of this additional latency, $\mathrm{T_{min}(B,L)}$, is given by
$\begin{split}\mathrm{T_{min}(B,L)}&=\mathrm{B}\cdot\left(\frac{\mathrm{M_{KVcache}}\cdot\mathrm{L}+\mathrm{M_{act}}(L)}{n_{\mathrm{acc}}\cdot\mathrm{BW_{mem}}}\right)+\mathrm{Comm(B,L)}\end{split}$
where $\mathrm{Comm(B,L)}$ denotes the communication overhead between accelerators. Increasing $B$ leads to larger KV\$ and activation sizes, thus increasing the minimum bound of TPOT. The theoretical maximum batch size, $\mathrm{B_{SLO}}$, that satisfies the SLO time limit ($\mathrm{TPOT_{SLO}}$) can be achieved under the minimum latency, and thus satisfies the following equation:
$\begin{split}\mathrm{TPOT_{SLO}}&=n_{\mathrm{dec}}\cdot\left(\frac{\mathrm{M_{attn}}\cdot{\mathrm{deg_{DP}}}+\mathrm{M_{MoE}}}{n_{\mathrm{acc}}\cdot\mathrm{BW_{Mem}}}\!+\!\mathrm{T_{min}(B_{SLO},L)}\right)\end{split}$
A batch size B greater than $\mathrm{B_{SLO}}$ can never satisfy the $\mathrm{TPOT_{SLO}}$ time limit, establishing an upper bound on the feasible batch size.
Observation-P6: While MoE tightens the batch size limit due to weight overhead, MLA complements this through its small KV\$ size.
MoE weights ($\mathrm{M_{MoE}}$) are typically larger than the FFN weights in standard LLMs. As a result, MoE increases memory requirements for the model weights, reducing $\mathrm{B_{cap}}$ as less memory space remains for KV\$ (see Eq. 11). It also increases model load latency, which shortens the time available for $\mathrm{T_{min}(B_{SLO},L)}$, thereby reducing $\mathrm{B_{SLO}}$. In contrast, the reduction of $\mathrm{M_{KVcache}}$ and $\mathrm{M_{attn}}$ by MLA enables storing KV\$ for more requests in main memory, thereby increasing $\mathrm{B_{cap}}$. It also reduces the load time for $\mathrm{M_{KVcache}}$ and $\mathrm{M_{attn}}$, allowing a higher $\mathrm{B_{SLO}}$ (see Eq. 12). Thus, as for batch size limits, MLA and MoE exhibit complementary effects.
VII End-to-End Model Execution Analysis
This section synthesizes our findings by analyzing the end-to-end execution of modern LLMs. We evaluate how the interplay between MLA, MoE, system scale, and interconnect bandwidth determines overall serving performance.
VII-A The Synergistic impact of MLA and MoE
LLMs using MLA and MoE achieve significantly higher throughput than conventional models. This is because MLA and MoE have a powerful synergistic relationship. MLA’s highly-compressed KV\$ dramatically increases the memory capacity available for batching ($\mathrm{B_{cap}}$). This, in turn, allows the system to form the large batches required to fully utilize the compute resources of the sparsely activated experts in MoE blocks, which would otherwise be constrained by memory.
Figure 9 illustrates this by comparing DeepSeek-R1 and GPT-3. For a sequence length of 8192, DeepSeek-R1’s $\mathrm{B_{cap}}$ (7360) is nearly 60$\times$ larger than GPT-3’s (124). Consequently, DeepSeek-R1 can be configured with a batch size large enough to approach its ridge point $\mathrm{B_{RP}}(=\!\mathrm{B_{attn}})$, whereas GPT-3 becomes memory-capacity-limited long before its compute resources can be saturated.
Observation-P7: MLA and MoE are synergistic. MLA’s KV\$ compression enables the large batch sizes, essential for MoE to achieve high a.i. and accelerator utilization.
Figure 10: System throughput and execution time ratio of operations in decode stage of DeepSeek-R1, varying batch sizes and sequence lengths. In this experiment, we assume a 64 B200 GPU system.
VII-B Scaling the system out
While the MLA/MoE combination is powerful, memory capacity can still become a bottleneck for very long sequences. A natural response is to scale the system out by adding more accelerators ($n_{\mathrm{acc}}$). However, this introduces a new trade-off. Increasing $n_{\mathrm{acc}}$ to raise the memory capacity limit ($\mathrm{B_{cap}}$) also increases the batch size required to saturate the attention blocks ($\mathrm{B_{attn}}$), as DP is distributed over more devices.
Figure 10 shows this effect when scaling from 32 to 64 GPUs. The increased memory capacity allows the system to reach the optimal batch size for MoE experts ($\mathrm{B_{MoE}}$). However, because $\mathrm{B_{attn}}$ also doubles, the system may still fall short of the batch size needed to saturate the attention layers, particularly for very long sequences where core attention latency remains a significant factor.
Observation-P8: Scaling out a system for longer sequences creates a trade-off. While adding accelerators increases memory capacity, it raises the batch size required to keep them compute-bound, requiring careful system balancing.
Figure 11: System throughput and execution time ratio of decode stage of DeepSeek-R1 when using InfiniBand XDR (800Gb/s) among a group of GPUs (DGX) varying batch sizes and sequence length. We assume 32 B200 GPU system.
VII-C The critical role of interconnect
The performance of a scaled-out MoE-based system is highly sensitive to interconnect bandwidth [10]. The all-to-all communication pattern, required to dispatch every token to its designated experts and then combine the results, creates dense network traffic that can easily become a bottleneck. As shown in Figure 11, moving from a high-bandwidth fabric like NVLink to a lower-bandwidth one like InfiniBand dramatically increases this communication overhead. Higher communication latency consumes a larger portion of the per-token time budget, which directly reduces the achievable batch size under a given SLO ($\mathrm{B_{SLO}}$) and leads to underutilization. Thus, using a high bisection bandwidth interconnect is critical for efficient system deployment.
Figure 12: Throughput comparison of 256 GPU and 32 GPU×$\times$8 systems of decode stage of DeepSeek-R1 when $L\!=\!2048$ and $L\!=\!16384$. 32GPUs are connected via 900 GB/s interconnect.
This sensitivity forces a critical deployment decision: using many small, tightly-coupled instances (e.g., 32 GPU×$\times$8) versus one large, monolithic instance (e.g., 256 GPU). Since it is difficult to scale the number of accelerators while maintaining high bisection bandwidth, we vary the interconnect bandwidth of the 256 GPU configuration to 900 GB/s, 300 GB/s, and 100 GB/s, which are equal to or lower than that of each 32 GPU instance.
As Figure 12 shows, the optimal choice depends on the workload. For shorter sequences, multiple small instances are more cost-effective because communication is contained within high-bandwidth domains, and the memory overhead of replicating MoE weights is manageable. When $L\!=\!2048$ (Figure 12(a)), at the batch size $\mathrm{B_{RP}}$ where throughput is maximized, 32 GPU×$\times$8 achieves equivalent throughput as 256 GPU with 900 GB/s interconnect bandwidth. At this point, in 256 GPU, each GPU is responsible for executing only one expert, but each expert processes 8 times more tokens than in 32 GPU×$\times$8. As the Op/B of experts in 256 GPU falls in a compute-bound region, it results in higher latency. Thus, the latency of MoE blocks becomes similar in both systems.
For very long sequences (e.g., $L\!=\!16384$ in Figure 12(b)), however, a single large instance is superior. The memory savings from storing the massive MoE weights only once by 256 GPU frees up system-wide capacity for a larger $\mathrm{B_{cap}}$, which is essential for handling the large KV\$, over 32 GPU×$\times$8. This leads to higher overall throughput, even if the large-scale interconnect has higher latency. For example, even with a reduced interconnect bandwidth of 300 GB/s, 256 GPU delivers better throughput by reducing MoE execution latency.
Observation-P9: High-bandwidth interconnects are critical for MoE performance. The optimal system topology depends on the sequence length, which dictates the trade-off between the communication overhead of a large monolithic system and the memory-capacity overhead of replicating weights across many smaller systems.
VIII Conclusion
Modern LLMs challenge the premise of specialized attention hardware. Architectural shifts to Multi-head Latent Attention (MLA) and Mixture-of-Experts (MoE) move the performance bottleneck away from memory-bound attention. With layer reordering, MLA becomes a nearly compute-bound task well-suited for contemporary accelerators, diminishing the case for dedicated hardware. The new design challenge is balancing the complementary demands of MLA and MoE. MoE requires large batches for efficiency, creating memory pressure from its large weights. MLA alleviates this by drastically reducing KV cache size, which in turn enables the large batches required for efficient MoE execution on long sequences. These findings demand a new system design methodology. Data parallelism proves more effective than tensor parallelism for reordered MLA. Further, the all-to-all communication in MoE makes high-bandwidth interconnects critical for meeting latency SLOs and maximizing throughput in large-scale systems.
Appendix A Symbol table
Table IV: Symbols used throughout this paper, their descriptions, and the exemplar parameters used in DeepSeek-R1 [11]
Term Description DeepSeek-R1    Term Description DeepSeek-R1
TP/DP/EP Tensor / Data / Expert Parallelism -    $\mathbf{O}_{t}$ Context output -
$\mathrm{deg_{TP}/_{DP}/_{EP}}$ TP / DP / EP degree -    $\mathbf{U}_{t}$ Final attention output / FFN input -
B Batch size -    $\mathbf{H}_{t}$ FFN output / Decoder block input -
$L$ Sequence length -    $\mathrm{RP_{device}}$ Ridge point of device -
$n_{\mathrm{dec}}$ Decoder blocks 61    Q, K, V Query, Key, Value -
$d_{\mathrm{emb}}$ Embedding dimension 7168    $\textbf{W}_{Q}$, $\mathbf{W}_{K}$, $\mathbf{W}_{V}$ Weight for Q, K, V generation -
$n_{\mathrm{hd}}$ Number of heads 128    $\mathbf{W}_{\mathrm{attn\_out}}$ Out projection weight in attention (16384, 7168)
$d_{\mathrm{hd}}$ Head dimension 128    $\mathbf{W}_{\mathrm{gate}}$, $\mathbf{W}_{\mathrm{up}}$ Weight for gate/up in FFN (7168, 18432)
$d_{\mathrm{dec}}$ Decompressed Q, KV dimension 16384    $\mathbf{W}_{\mathrm{down}}$ Weight for down in FFN (18432, 7168)
$d_{\mathrm{Qco}}$, $d_{\mathrm{KVco}}$ Compressed Q, KV dimension 1536, 512    $\mathbf{W}_{\mathrm{route}}$ MoE route weight (7168, 256)
$d_{\mathrm{RoPE}}$ Rotary PE dimension 64    $\mathbf{W}_{\mathrm{exp_{n},\,gate}}$, $\mathbf{W}_{\mathrm{exp_{n},\,up}}$, $\mathbf{W}_{\mathrm{exp_{n},\,down}}$ MoE up/down projection weights (7168, 2048), (7168, 2048), (2048, 7168)
$d_{\mathrm{MoE}}$ MoE dimension 2048    $\mathbf{W}_{\mathrm{CQ}}$, $\mathbf{W}_{\mathrm{CKV}}$ Q comp / KV compression (7168, 1536), (7168, 512)
$n_{\mathrm{k}}$ Top-k experts 8    $\mathbf{W}_{\mathrm{DQ}}$ Q decompression weight (1536, 16384)
$n_{\mathrm{e}}$ Number of experts 256    $\mathbf{W}_{\mathrm{DK}}$, $\mathbf{W}_{\mathrm{DV}}$ K, V decompression weights (512, 16384)
$\mathbf{Q}_{\mathrm{NoPE}}$ Query vector (No RoPE) (1, 16384)    $\mathbf{W}_{\mathrm{RQ}}$ RoPE Q weight (1536, 8192)
$\mathbf{Q}_{\mathrm{RoPE}}$ Query after RoPE (1, 8192)    $\mathbf{W}_{\mathrm{RK}}$ RoPE K weight (7168, 64)
$\mathbf{K}_{\mathrm{RoPE}}$ Key vector for RoPE (1, 64)    $\mathbf{C}_{Q}$ Latent Q (compressed) -
$\mathbf{S}_{\mathrm{RoPE}}$ Score output with RoPE -    $\mathbf{C}_{\mathrm{KV}}$ Latent KV (compressed) -
$\mathbf{S}_{\mathrm{NoPE}}$ Score output without RoPE -    $n_{\mathrm{acc}}$ Number of accelerators -
$\mathbf{S}_{t}$ Final score output -    - -